# Robustness of deep learning classification to adversarial input on GPUs: asynchronous parallel accumulation is a source of vulnerability

### v1

@@ 5 Discussion and Conclusions
We show APFPR has significant impacts on classification accuracy and robustness. We developed a novel black-box Bayesian optimization attack (EWA) to determine properties of additional workloads that reliably result in misclassification (up to 4.6% accuracy decrease, at least 75% of the time). While our current work focused on matrix multiplications as the external background workload, future research will explore more complex workloads in both isolated and cloud environments. Additionally, GPU scheduler probing tools need to be developed to investigate how the EWA and associated workloads impact misclassification.
We introduced the LP approach to efficiently identify permutations that maximize prediction errors, offering significant advantages over brute force search. Our results demonstrate that both run-to-run variability and EWA can be bound by the LP worst-case estimates. Direction for future work involve optimizing LP further, integrating it into existing robustness verification tools and performing more exhaustive testing over different ML architectures and datasets.
Our examination of GPU system states—including varying workloads, partitions, and power settings—showed significant influence on the ordering of parallel operations across three different GPU models (from NVIDIA and AMD). These findings highlight that testing with a single GPU configuration is insufficient to fully account for non-determinism in model performance. This further emphasizes the value of our LP approach over repeated inferences, which would otherwise require extensive testing across multiple GPU and GPU states.
@@ 6 Hardware and Systems Used in Experiments
Tests on GH200 GPUs are run on two separate compute nodes, one running SLE 15 (enterprise) and the other Red Hat Enterprise Linux 9.4 (Plow) with 2 NVIDIA GH200 GPUs and 72-core ARM Neoverse-V2 CPUs. V100 tests use an IBM Power System AC922 (Red Hat 8) with dual Power9 CPUs and six V100 GPUs (16GB HBM2). MI250X tests are performed on an HPE Cray EX (SLE 15) with AMD EPYC CPUs and four MI250X GPUs (eight GCDs per node). H100 tests run on Ubuntu 22.04.06 with two 40GB H100 GPUs and an AMD EPYC 7302 CPU. We use PyTorch 2.4, PyTorch Geometric 2.6 and CUDA 12.0.

### v2

@@ 5 Discussion and Conclusions
We show that APFPR has significant impacts on classification accuracy and robustness. We developed a novel black-box Bayesian optimization attack (EWA) to determine the properties of additional workloads that reliably result in misclassification (up to 4.6% accuracy decrease, at least 75% of the time). While our current work focused on matrix multiplications as the external background workload, future research should explore more complex workloads in both isolated and cloud environments. Additionally, GPU scheduler probing tools must be developed to investigate how EWA and associated workloads impact misclassification. We introduced the LP approach to efficiently identify permutations that maximize prediction errors, offering significant advantages over brute-force search. Our results demonstrate that both run-to-run variability and EWA can be bound by the LP worst-case estimates. Direction for future work involves optimizing LP further, integrating it into existing robustness verification tools and performing more exhaustive testing over different ML architectures and datasets.
Our examination of GPU system states, including varying workloads, partitions, and power settings, showed a significant influence on the ordering of parallel operations across three different GPU models (from NVIDIA and AMD). These findings highlight that testing with a single GPU type is insufficient to fully account for non-determinism in model performance. This further emphasizes the value of our LP approach over repeated inferences, which would otherwise require extensive testing across multiple GPUs and GPU states. While frameworks like PyTorch can support deterministic operations, non-deterministic kernels and atomic operations are deeply integrated, making full determinism costly to implement. Workarounds such as integer quantization may help but often reduce accuracy, particularly in deep architectures. Purpose-built deterministic hardware such as the Groq LPU offers a valuable alternative benchmark for reliable inference.
@@ 6 Hardware and Systems Used in Experiments
Tests for the V100 are run on the Summit supercomputer at the Oak Ridge Leadership Computing Facility (OLCF), running Redhat OS 8. Summit is an IBM system; each IBM Power System AC922 node has two Power9 CPUs with 512 GB of memory and 6 V100 NVIDIA GPU with 16GB of HBM2 memory.
Tests on Mi250X AMD GPU are obtained on the Frontier supercomputer at OLCF, running SLE 15 (enterprise). Frontier is an HPE Cray EX supercomputer; each Frontier compute node has a 64-core AMD “Optimized 3rd Gen EPYC” CPU with 512 GB of DDR4 memory and 4 AMD MI250X GPUs, each with 2 Graphics Compute Dies (GCDs) for a total of 8 GCDs per node.
Tests on GH200 GPUs are run on two separate compute nodes, one running SLE 15 (enterprise) and the other Red Hat Enterprise Linux 9.4 (Plow) with 2 NVIDIA GH200 GPUs and 72-core ARM Neoverse-V2 CPUs. H100 tests run on Ubuntu 22.04.06 with two 40GB H100 GPUs and an AMD EPYC 7302 CPU. We use PyTorch 2.4, PyTorch Geometric 2.6 and CUDA 12.0.
Acknowledgments
This work was supported in part by the ORNL AI LDRD Initiative, the Swiss Platform For Advanced Scientific Computing (PASC), and the Accelerated Data Analytics and Computing Institute (ADAC). It used resources of the OLCF, a DOE Office of Science User Facility [DE-AC05-00OR22725], and the Swiss National Supercomputing Centre. The authors thank Hayashi Akihiro and Pim Witlox for insightful discussions.
Disclosure of Interests
The authors have no competing interests to declare that are relevant to the content of the article.